# Improving Speech Emotion Recognition with Mutual Information Regularized Generative Model

### v2

@@ Abstract
Data scarcity remains a fundamental challenge in speech emotion recognition (SER), particularly as models grow more complex and the demand for multimodal systems increases. While generative data augmentation offers a promising solution, existing approaches often produce emotionally inconsistent samples due to oversimplified conditioning on categorical labels. This paper introduces a novel mutual-information-regularised generative framework that addresses these limitations through cross-modal information transfer. Our method leverages textual cues to guide the synthesis of emotionally relevant audio features while employing mutual information as a quantitative regulariser to ensure strong dependency between generated samples and their conditioning signals. We extend this approach to multimodal settings, enabling the generation of novel, paired (audio, text) samples. Comprehensive evaluation on three benchmark datasets (IEMOCAP, MSP-IMPROV, MSP-Podcast) demonstrates that our framework consistently outperforms existing augmentation methods, achieving state-of-the-art performance with improvements of up to 2.6% in unimodal SER and 3.2% in multimodal emotion recognition. Most importantly, we demonstrate that mutual information functions as both a regulariser and a measurable metric for generative quality, offering a systematic approach to data augmentation in affective computing.
@@ 1 Introduction
Speech Emotion Recognition (SER) has advanced considerably over the past decade, driven by the success of deep learning architectures that can capture the complex acoustic and linguistic cues underlying human emotions. These developments have led to tremendous improvements in classification accuracy across widely used benchmark datasets [1, 2]. Nevertheless, compared with computer vision or natural language processing, SER research continues to face a major bottleneck: the lack of large, high-quality, and consistently annotated emotional speech corpora. This scarcity of data constrains model generalisation and makes data augmentation a crucial strategy for building reliable SER systems.
@@
Conventional augmentation techniques, such as time-stretching, pitch-shifting, or adding Gaussian noise, expand data volume but essentially produce distorted copies of existing recordings [3, 4]. They rarely introduce genuinely new emotional variation. In contrast, generative models can synthesise novel data by learning the underlying distribution of emotional speech. Among these, Generative Adversarial Networks (GANs) [5] have gained particular traction due to their ability to produce realistic, high-fidelity features, and several studies have reported improved SER performance through GAN-based augmentation [6, 2].
@@
Despite their promise, most current generative approaches depend on class-conditioned GANs that assume a deterministic mapping between emotion labels and generated samples. This assumption oversimplifies human emotion, which is inherently complex and context-dependent. As a result, generated data may lack the diversity or authenticity needed for robust emotion modelling. To mitigate this issue, recent work in representation learning has explored mutual-information-based regularisation, which encourages stronger and more interpretable relationships between latent variables and generated outputs [7, 8]. By maximising mutual information, it can be quantitatively ensured that the generated features retain meaningful emotional structure, an essential property for effective augmentation.
@@
At the same time, emotion recognition is increasingly being treated as a multimodal task. Text and audio provide different but complementary emotional signals. While the words we choose express the meaning behind our feelings, the tone of our voice, like prosody and intonation adds a layer of expressive nuance. This duality helps convey emotions more richly and effectively [9, 10]. Although multimodal fusion has become common, only a few studies have explored multimodal augmentation itself. Leveraging text information to guide audio generation could substantially improve emotion modelling, yet this area remains under-explored.
@@
Motivated by these gaps, this study introduces a mutual-information-regularised generative augmentation framework for SER. Building on the InfoGAN architecture [7], our approach synthesises emotion-aware speech features while preserving strong dependencies between generated samples, emotion labels, and text embeddings. We further extend this concept to a multimodal setting by combining both audio and text representations, enabling the generation of new paired (audio, text) features for training emotion classifiers. The key contributions of this work are:
@@ 2.1 Generative Data Augmentation for SER
The limited availability of high-quality, well-annotated emotional speech data has made data augmentation a critical component of modern SER. Traditional approaches like time-stretching, pitch-shifting, and adding noise can increase the quantity of data, but they often fail to capture the rich emotional variations necessary for accurate recognition [4, 11].
To tackle these shortcomings, researchers have utilised deep generative models that can learn the complex variations in emotional speech. Generative Adversarial Networks (GANs) [12] are popular in this space because they can generate high-quality features and improve the robustness of models [13, 6]. For example, Latif et al. [13] have introduced adversarial autoencoders specifically for SER, while Yi et al. [6] built upon this with specialised generators. More recently, Latif et al. [14] combined GANs with mix-up strategies, and Wang et al. [15] incorporated triplet loss to tackle issues with class imbalance.
Sahu et al [16] introduced mutual information maximisation through InfoGAN [7] to enhance emotion-controllable synthesis. Building on this recent studies have explored new techniques: Latif  et al. [17] worked on emotional text-to-speech generation, Malik et al. [18] applied diffusion models for mel-spectrogram synthesis, and Kim et al. [19] combined VAEs and diffusion techniques to boost data realism.
Further pushing the envelope, Naji et al. [20] introduced a transformer-based diffusion framework that achieves near-human-level generation quality for emotional speech. Meanwhile, Roh and Lee [21] brought forth a conditional DCGAN to better control the emotional intensity in generated samples. Shilandari et al. [22] also explored emotion-style transfer with CycleGANs, resulting in a noticeable improvement in emotional diversity.
Limitations and Gap: Despite the advancements, many generative methods simplify the relationship between emotion labels and the data they generate, overlooking the complexities of emotional variability. Few studies explicitly connect the latent and generated representations. Although some use mutual information or textual conditioning [16, 18], there is a lack of systematic mutual-information regularisation that could enhance the emotional coherence between generated audio and cross-modal signals.
@@ 2.2 Multimodal and Cross-Modal Learning
Recognition of emotions is inherently a multimodal endeavor: textual semantics and acoustic prosody offer complementary insights [10]. Advances in multimodal deep learning have fostered the ability to co-learn across modalities, which improves both robustness and generalisation [23].
Typically, conventional methods focus on fusion or alignment of modalities [24, 25], while cross-modal transfer leverages one modality to inform another. Recent research has utilised contrastive learning to strengthen shared representations [26] and attention-based fusion to dynamically weigh contributions from each modality [24].
@@ 2.3 Handling Missing Modalities
Multimodal systems frequently contend with partial or noisy inputs. Researchers like Gonçalves et al. [24] have employed multitask learning to maintain robustness when certain modalities are missing, while Wang et al. [27] explored random masking techniques. Meng et al. [28] have even used autoencoders to reconstruct incomplete signals. More recent work by Wang et al. [29] has continued to push this research agenda forward, refining methods for better recovery and analysis in multimodal scenarios.
@@ 2.4 Our Contribution and Positioning
Latif et al. [13] 2020
Yi et al. [6] 2020
Latif et al. [14] 2020
Sahu et al. [16] 2022
Wang et al. [15] 2022
Latif et al. [17] 2023
Malik et al. Model [18] 2023
Kim et al. [19] 2024
Goncalves et al. [24] 2022
Chen et al. [25] 2023
Wang et al. [27] 2023
Meng et al. [28] 2023
Wang et al. [29] 2023
Liu et al. [26] 2024
@@ 3 Proposed Framework
Figure 1: The visual summary of proposed augmentation framework to improve SER, consisted of three stages. First stage (top left): baseline model is trained with contrastive loss and InfoNCE loss. Second stage (top right): InfoGAN, GAN with mutual information module that predicts latent $c^{\prime}$ and $t^{\prime}$ from generated $\hat{h}$, is trained and mutual information module re-uses the same prediction layers in the first stage. Third stage (bottom): we have two parallel stream in the final stage, SER case (bottom left) and multimodal SER case (bottom right), which is training the linear classification module with all possible input combinations by switching between $h$ or $\hat{h}$ and $t$ or $t^{\prime}$. linear classifier layer is fine tuned with generated samples.
@@ 3.1 Baseline Model with Cross-Modal Alignment
The first stage establishes a robust feature space. We fine-tune a pre-trained audio transformer (e.g., AST [30] or Wav2Vec2 [31]) on the target SER task. The model takes a raw audio input $x_{a}$ and produces an audio feature vector:
@@
To inject cross-modal structure and prepare for multimodal augmentation, we align the audio features with textual embeddings. Using the ground-truth transcript $x_{t}$, a frozen text transformer (e.g., BERT [32]) produces a text feature:
@@
Contrastive Alignment Loss ($\mathcal{L}_{CL}$): Inspired by CLAP [33], this loss pulls corresponding $(h^{i},t^{i})$ pairs together while pushing apart non-matching pairs within a mini-batch:
@@
Mutual Information Loss ($\mathcal{L}_{MI}$): To further maximise the dependency between $h$ and $t$, we employ an InfoNCE loss [34]. We project $h$ into the text embedding space, $\hat{t}=W_{t}h+b_{t}$, and compute:
@@ 3.2 Mutual Information Regularised Feature Generation
The second stage learns a generative model to synthesize new audio features from the distribution of the aligned feature space. We build upon the InfoGAN architecture [7], which augments the standard GAN objective with a mutual information term to ensure the generated samples are interpretable and controllable.
@@ 3.3 Data Augmentation and Classifier Fine-tuning
This staged approach ensures that the augmented data consists of novel, emotionally coherent, and cross-modally consistent samples that directly address the core challenges of data scarcity in SER.
@@ 4.2.1 Data augmentation for SER
We adopt the pre-trained AST [30] from ‘https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593’ as backbone. And add linear classifier after pooled output from AST. For text feature encoder, pre-trained BERT [32], that is ’bert-uncased-base’ version. For SER experiment, we do not consider multimodal SER accuracy, thus, BERT is frozen and we only use it to extract text features from ground-truth transcript.
We train InfoGAN to generate feature encoding, which is in the dimension of 768. So we take the simplest form of architecture for Discriminator and Generator which is a linear layer without any activation function. Also the network for mutual information maximization is also linear layer only. Especially, the linear layer that maximizes mutual information with $y$ uses the same layer from emotion classifier. In this way, we can prevent latent code vector to encode something other than emotion, as there are many other complex variations inherent in the training dataset. Emotional code vector is trained with cross entropy to maximize the mutual information. For the latent embedding given from BERT embedding we use InfoNCE loss, which is a contrastive learning loss trying attract the BERT embedding and predicted text embedding from generated sample. Lastly, we adopted mix-up strategy for training generator and discriminator as proposed in [4], for stable training of generator.
@@ 4.2.2 Data augmentation for multimodal SER
We adopt the pre-trained Wav2Vec2 [36] from ‘https://huggingface.co/audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim’ for multimodal SER experiment. For text feature encoder, pre-trained RoBERTA [32], that is ’roberta-uncased-large’ version. For multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage. After that RoBERTa model weight is frozen during training baseline and onwards. For MSP-Podcast experiment, we use Llama 2.0 [37] from ‘https://huggingface.co/meta-llama/Llama-2-7b’ instead of RoBERTa for text features, but same Wav2Vec2 for audio features. Other than the change of feature size, due to change of back-bone model, configuration for InfoGAN stay the same. Multimodal SER classification layer takes concatenated feature of both modalities. The classifier consists linear layer, one hidden layer (256) or two hidden layers (256-128), for IEMOCAP, MSP-IMPROV or MSP-Podcast respectively.
@@ 4.2.3 Hyperparameters
We used Adam optimizer with 1e41e-4 learning rate for training. We employed different weights per class during cross entropy loss computation due to class imbalance, for MSP-IMPROV and MSP-Podcast. The weights were inverse proportional to the total numbers of samples per each class. Lora [38] was adopted when Llama was trained for text emotion recognition. The weighting factors for each loss terms ($\alpha$, $\beta$, $\lambda_{y}$ and $\lambda_{t}$) were fixed to $1$.
@@ 5 Result
The experiment was conducted in two rounds. First, our data augmentation framework was tested on SER case. In first round, we observe the performance difference before and after data augmentation, and ablation study to check the contributing effects of each modules in the model. Finally, we run the multimodal SER experiments. The second round, we compare the performance of the model using proposed framework with existing works.

### v3

@@ Abstract
Lack of large, well- annotated emotional speech corpora continues to limit the performance and robustness of speech emotion recognition (SER), particularly as models grow more complex and the demand for multimodal systems increases. While generative data augmentation offers a promising solution, existing approaches often produce emotionally inconsistent samples due to oversimplified conditioning on categorical labels. This paper introduces a novel mutual-information-regularised generative framework that combines cross-modal alignment with feature-level synthesis. Building on an InfoGAN-style architecture, our method first learns a semantically aligned audio–text representation space using pre-trained transformers and contrastive objectives. A feature generator is then trained to produce emotion-aware audio features while employing mutual information as a quantitative regulariser to ensure strong dependency between generated features and their conditioning variables. We extend this approach to multimodal settings, enabling the generation of novel, paired (audio, text) features. Comprehensive evaluation on three benchmark datasets (IEMOCAP, MSP-IMPROV, MSP-Podcast) demonstrates that our framework consistently outperforms existing augmentation methods, achieving state-of-the-art performance with improvements of up to 2.6% in unimodal SER and 3.2% in multimodal emotion recognition. Most importantly, we demonstrate that mutual information functions as both a regulariser and a measurable metric for generative quality, offering a systematic approach to data augmentation in affective computing.
@@ 1 Introduction
Speech Emotion Recognition (SER) has advanced considerably over the past decade, driven by the success of deep learning architectures that can capture the complex acoustic and linguistic cues underlying human emotions. These developments have led to tremendous improvements in classification accuracy across widely used benchmark datasets [Schuller2018SER, Latif2021SurveySER]. Nevertheless, compared with computer vision or natural language processing, SER research continues to face a major bottleneck: the lack of large, high-quality, and consistently annotated emotional speech corpora. This scarcity of data constrains model generalisation and makes data augmentation a crucial strategy for building reliable SER systems.
@@
Conventional augmentation techniques, such as time-stretching, pitch-shifting, or adding Gaussian noise, expand data volume but essentially produce distorted copies of existing recordings [Parthasarathy2020Ladder, zhang2017mixup]. They rarely introduce genuinely new emotional variation. In contrast, generative models can synthesise novel data by learning the underlying distribution of emotional speech. Among these, Generative Adversarial Networks (GANs) [Goodfellow2014GAN] have gained particular traction due to their ability to produce realistic, high-fidelity features, and several studies have reported improved SER performance through GAN-based augmentation [Yi2020AdversarialSER, Latif2021SurveySER].
@@
Despite their promise, most current generative approaches depend on class-conditioned GANs that assume a deterministic mapping between emotion labels and generated samples. This assumption oversimplifies human emotion, which is inherently complex and context-dependent. As a result, generated data may lack the diversity or authenticity needed for robust emotion modelling. To mitigate this issue, recent work in representation learning has explored mutual-information-based regularisation, which encourages stronger and more interpretable relationships between latent variables and generated outputs [chen2016infogan, Belghazi2018MINE]. By maximising mutual information, it can be quantitatively ensured that the generated features retain meaningful emotional structure, an essential property for effective augmentation.
@@
At the same time, emotion recognition is increasingly being treated as a multimodal task. Text and audio provide different but complementary emotional signals. While the words we choose express the meaning behind our feelings, the tone of our voice, like prosody and intonation adds a layer of expressive nuance. This duality helps convey emotions more richly and effectively [Poria2017ContextSentiment, Busso2019MSPPodcast]. Although multimodal fusion has become common, only a few studies have explored multimodal augmentation itself. Leveraging text information to guide audio generation could substantially improve emotion modelling, yet this area remains under-explored.
@@
Motivated by these gaps, this study introduces a mutual-information-regularised generative augmentation framework for SER. Building on the InfoGAN architecture [chen2016infogan], our approach synthesises emotion-aware speech features while preserving strong dependencies between generated features, emotion labels, and text embeddings. We further extend this concept to a multimodal setting by combining both audio and text representations, enabling the generation of new paired (audio, text) features for training emotion classifiers. The key contributions of this work are:
@@ 2.1 Generative Data Augmentation for SER
The limited availability of high-quality, well-annotated emotional speech data has made data augmentation a critical component of modern SER. Traditional approaches like time-stretching, pitch-shifting, and adding noise can increase the quantity of data, but they often fail to capture the rich emotional variations necessary for accurate recognition [zhang2017mixup, parthasarathy2020semi].
To tackle these shortcomings, researchers have utilised deep generative models that can learn the complex variations in emotional speech. Generative Adversarial Networks (GANs) [goodfellow2014generative] are popular in this space because they can generate high-quality features and improve the robustness of models [latif2020multi, Yi2020AdversarialSER]. For example, Latif et al. [latif2020multi] have introduced adversarial autoencoders specifically for SER, while Yi et al. [Yi2020AdversarialSER] built upon this with specialised generators. More recently, Latif et al. [latif2020augmenting] combined GANs with mix-up strategies, and Wang et al. [wang2022generative] incorporated triplet loss to tackle issues with class imbalance.
Sahu et al [sahu2020modeling] introduced mutual information maximisation through InfoGAN [chen2016infogan] to enhance emotion-controllable synthesis. Building on this recent studies have explored new techniques: Latif  et al. [latif2023generative] worked on emotional text-to-speech generation, Malik et al. [malik2023preliminary] applied diffusion models for mel-spectrogram synthesis, and Kim et al. [kim2024generation] combined VAEs and diffusion techniques to boost data realism.
Further pushing the envelope, Naji et al. [naji2025transformerSER] introduced a transformer-based diffusion framework that achieves near-human-level generation quality for emotional speech. Meanwhile, Roh and Lee [roh2024conditionalDCGAN] brought forth a conditional DCGAN to better control the emotional intensity in generated samples. Shilandari et al. [shilandari2024cycleganSER] also explored emotion-style transfer with CycleGANs, resulting in a noticeable improvement in emotional diversity.
Limitations and Gap: Despite the advancements, many generative methods simplify the relationship between emotion labels and the data they generate, overlooking the complexities of emotional variability. Few studies explicitly connect the latent and generated representations. Although some use mutual information or textual conditioning [sahu2020modeling, malik2023preliminary], there is a lack of systematic mutual-information regularisation that could enhance the emotional coherence between generated audio and cross-modal signals.
@@ 2.2 Multimodal and Cross-Modal Learning
Recognition of emotions is inherently a multimodal endeavor: textual semantics and acoustic prosody offer complementary insights [Busso2019MSPPodcast]. Advances in multimodal deep learning have fostered the ability to co-learn across modalities, which improves both robustness and generalisation [liang2024foundations].
Typically, conventional methods focus on fusion or alignment of modalities [goncalves2022robust, chen2023inter], while cross-modal transfer leverages one modality to inform another. Recent research has utilised contrastive learning to strengthen shared representations [liu2024contrastive] and attention-based fusion to dynamically weigh contributions from each modality [goncalves2022robust].
@@ 2.3 Handling Missing Modalities
Multimodal systems frequently contend with partial or noisy inputs. Researchers like Gonçalves et al. [goncalves2022robust] have employed multitask learning to maintain robustness when certain modalities are missing, while Wang et al. [wang2023exploring] explored random masking techniques. Meng et al. [meng2024deep] have even used autoencoders to reconstruct incomplete signals. More recent work by Wang et al. [wang2024incomplete] has continued to push this research agenda forward, refining methods for better recovery and analysis in multimodal scenarios.
@@ 2.4 Our Contribution and Positioning
Latif et al. [latif2020multi] 2020
Yi et al. [Yi2020AdversarialSER] 2020
Latif et al. [latif2020augmenting] 2020
Sahu et al. [sahu2020modeling] 2022
Wang et al. [wang2022generative] 2022
Latif et al. [latif2023generative] 2023
Malik et al. Model [malik2023preliminary] 2023
Kim et al. [kim2024generation] 2024
Goncalves et al. [goncalves2022robust] 2022
Chen et al. [chen2023inter] 2023
Wang et al. [wang2023exploring] 2023
Meng et al. [meng2024deep] 2023
Wang et al. [wang2024incomplete] 2023
Liu et al. [liu2024contrastive] 2024
@@ 3 Proposed Framework
Figure 1: The visual summary of proposed augmentation framework to improve SER, consisted of three stages. First stage (top left): baseline model is trained with contrastive loss and InfoNCE loss. Second stage (top right): InfoGAN, GAN with mutual information module that predicts latent $c^{\prime}$ and $t^{\prime}$ from generated $\hat{h}$, is trained and mutual information module re-uses the same prediction layers in the first stage. Third stage (bottom): we have two parallel stream in the final stage, SER case (bottom left) and multimodal SER case (bottom right), which is training the linear classification module with all possible input combinations by switching between $h$ or $\hat{h}$ and $t$ or $t^{\prime}$. linear classifier layer is fine tuned with generated features.
@@ 3.1 Baseline Model with Cross-Modal Alignment
The first stage establishes a robust feature space. We fine-tune a pre-trained audio transformer (e.g., AST [gong21b_interspeech] or Wav2Vec2 [baevski2020wav2vec]) on the target SER task. The model takes a raw audio input $x_{a}$ and produces an audio feature vector:
@@
To inject cross-modal structure and prepare for multimodal augmentation, we align the audio features with textual embeddings. Using the ground-truth transcript $x_{t}$, a frozen text transformer (e.g., BERT [devlin2018bert]) produces a text feature:
@@
Contrastive Alignment Loss ($\mathcal{L}_{CL}$): Inspired by CLAP [elizalde2023clap], this loss pulls corresponding $(h^{i},t^{i})$ pairs together while pushing apart non-matching pairs within a mini-batch:
@@
Mutual Information Loss ($\mathcal{L}_{MI}$): To further maximise the dependency between $h$ and $t$, we employ an InfoNCE loss [oord2018representation]. We project $h$ into the text embedding space, $\hat{t}=W_{t}h+b_{t}$, and compute:
@@ 3.2 Mutual Information Regularised Feature Generation
The second stage learns a generative model to synthesize new audio features from the distribution of the aligned feature space. We build upon the InfoGAN architecture [chen2016infogan], which augments the standard GAN objective with a mutual information term to ensure the generated features are interpretable and controllable.
@@ 3.3 Data Augmentation and Classifier Fine-tuning
This staged approach ensures that the augmented data consists of novel, emotionally coherent, and cross-modally consistent features that directly address the core challenges of data scarcity in SER.
@@ 4.2.1 Data augmentation for SER
We adopt the pre-trained AST [gong21b_interspeech] from ‘https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593’ as backbone. And add linear classifier after pooled output from AST. For text feature encoder, pre-trained BERT [devlin2018bert], that is ’bert-uncased-base’ version. For SER experiment, we do not consider multimodal SER accuracy, thus, BERT is frozen and we only use it to extract text features from ground-truth transcript.
We train InfoGAN to generate feature encoding, which is in the dimension of 768. So we take the simplest form of architecture for Discriminator and Generator which is a linear layer without any activation function. Also the network for mutual information maximization is also linear layer only. Especially, the linear layer that maximizes mutual information with $y$ uses the same layer from emotion classifier. In this way, we can prevent latent code vector to encode something other than emotion, as there are many other complex variations inherent in the training dataset. Emotional code vector is trained with cross entropy to maximize the mutual information. For the latent embedding given from BERT embedding we use InfoNCE loss, which is a contrastive learning loss trying attract the BERT embedding and predicted text embedding from generated sample. Lastly, we adopted mix-up strategy for training generator and discriminator as proposed in [zhang2017mixup], for stable training of generator.
@@ 4.2.2 Data augmentation for multimodal SER
We adopt the pre-trained Wav2Vec2 [10089511] from ‘https://huggingface.co/audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim’ for multimodal SER experiment. For text feature encoder, pre-trained RoBERTA [devlin2018bert], that is ’roberta-uncased-large’ version. For multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage. After that RoBERTa model weight is frozen during training baseline and onwards. For MSP-Podcast experiment, we use Llama 2.0 [touvron2023llama] from ‘https://huggingface.co/meta-llama/Llama-2-7b’ instead of RoBERTa for text features, but same Wav2Vec2 for audio features. Other than the change of feature size, due to change of back-bone model, configuration for InfoGAN stay the same. Multimodal SER classification layer takes concatenated feature of both modalities. The classifier consists linear layer, one hidden layer (256) or two hidden layers (256-128), for IEMOCAP, MSP-IMPROV or MSP-Podcast respectively.
@@ 4.2.3 Hyperparameters
We used Adam optimizer with 1e41e-4 learning rate for training. We employed different weights per class during cross entropy loss computation due to class imbalance, for MSP-IMPROV and MSP-Podcast. The weights were inverse proportional to the total numbers of samples per each class. Lora [hu2022lora] was adopted when Llama was trained for text emotion recognition. The weighting factors for each loss terms ($\alpha$, $\beta$, $\lambda_{y}$ and $\lambda_{t}$) were fixed to $1$.
@@ 5 Result
The experiment was conducted in two rounds. First, our data augmentation framework was tested on SER case. We observe the performance difference before and after data augmentation, and ablation study to check the contributing effects of each modules in the model. Finally, we run the multimodal SER experiments.